# Multi-lingual Functional Evaluation for Large Language Models

### v1

@@ B.3 High to X Low-Resource Language Gaps
Aya-23-35B 33.09 57.4 54.12
Mixtral-8x7B-Instruct 36.41 56.8 57.11
Mistral-7B-Instruct 33.28 44.2 50.11
Aya-expanse-32B 56.38 71.6 60.45
Gemma-2-9b-it 53.42 64.4 52.44
Qwen3-8b 67.65 73.0 62.89
@@ Appendix D Full Functional Benchmark Results
Aya-23-35B 46.21 31.60 30.31 35.86 36.23 13.12
Mixtral-8x7B-Instruct 45.84 31.42 29.21 25.32 19.78 9.43
Mistral-7B-Instruct 46.40 34.01 35.30 19.96 17.92 13.12
Aya-expanse-32B 70.24 50.83 46.77 44.18 41.40 13.86
Gemma-2-9b-it 69.87 51.94 48.42 45.1 42.69 16.45
Qwen3-8b 81.88 57.67 54.52 48.79 47.31 14.23
@@
English GPT-4o-mini 81.14 86.81 84.47 89.32
AYA 23-35B 46.21 56.83 49.16 60.07
AYA Exp-32B 70.24 77.94 78.74 84.89
Claude Sonnet 85.76 90.04 89.64 92.92
Mistral 46.40 57.79 57.12 67.87
Mixtral 45.84 57.19 55.08 65.95
Gemma2-9B 69.87 78.77 73.75 82.01
Qwen3-8B 81.88 88.00 84.84 89.92
French GPT-4o-mini 55.08 66.42 58.22 69.42
AYA 23-35B 31.60 43.16 34.75 46.76
AYA Exp-32B 50.83 61.87 57.86 68.35
Claude Sonnet 62.84 72.42 63.58 73.50
Mistral 34.01 45.92 43.99 55.88
Mixtral 31.42 43.65 37.71 51.08
Gemma2-9B 51.94 62.94 53.41 64.74
Qwen3-8B 57.67 69.06 59.51 70.50
Spanish GPT-4o-mini 53.23 64.62 56.74 68.10
AYA 23-35B 30.31 42.20 33.82 45.56
Gemma 7B 30.13 41.13 34.57 46.88
AYA Exp-32B 46.77 58.51 53.23 65.35
Claude Sonnet 58.96 69.42 60.25 70.62
Mistral 35.30 46.16 37.70 48.92
Mixtral 29.21 40.53 36.41 49.28
Gemma2-9B 48.42 60.55 51.20 63.18
Qwen3-8B 54.52 66.30 56.56 68.22
Yoruba GPT-4o-mini 35.67 47.72 39.00 51.31
AYA 23-35B 13.12 23.50 14.23 24.58
AYA Exp-32B 13.86 23.86 20.33 32.61
Claude Sonnet 34.19 46.28 37.89 50.59
Mistral 13.12 24.22 17.56 30.94
Mixtral 9.43 19.06 13.86 25.90
Gemma2-9B 16.45 29.01 17.92 31.29
Qwen3-8B 14.23 25.17 15.71 27.21
Arabic GPT-4o-mini 47.87 61.15 50.27 63.42
AYA 23-35B 35.86 50.72 39.74 54.68
AYA Exp-32B 44.18 58.27 47.87 61.75
Claude Sonnet 53.23 65.10 54.15 65.94
Mistral 19.96 33.69 26.06 40.17
Mixtral 25.32 38.25 37.79 46.16
Gemma2-9B 45.10 57.79 46.39 59.35
Qwen3-8B 48.79 61.39 49.53 62.11
Hindi GPT-4o-mini 50.27 62.11 51.94 63.78
AYA 23-35B 36.23 47.48 40.48 53.12
AYA Exp-32B 41.40 53.12 46.95 59.11
Claude Sonnet 52.31 63.54 53.60 64.86
Mistral 17.92 29.25 24.21 37.77
Mixtral 19.78 31.18 26.25 38.97
Gemma2-9B 42.69 55.51 44.73 57.91
Qwen3-8B 47.31 59.35 49.53 61.75
@@ F.2 CL-GSMSym Plots
Appendix G Failure Case Examples for CL-IFEval
[…]
Appendix H Template Examples for CL-GSMSym
@@
Template 10 Example (Spanish) Language: Spanish Template: 10 Question: Cuando Emma observa a su primo, saca una variedad de juguetes para él. La bolsa de bloques de construcción tiene 74 bloques dentro. El contenedor de animales de peluche tiene 37 animales de peluche dentro. La torre de anillos apilables tiene 30 anillos multicolores. Emma compró recientemente un tubo de pelotas saltarinas, lo que eleva el número total de juguetes para su primo a 215. ¿Cuántas pelotas saltarinas vinieron en el tubo? Answer: Sea $T$ el número de pelotas que rebotan en el tubo. Después de comprar el tubo de pelotas, Emma tiene $21574+37+30+T=141+T=$ 215 juguetes. Por lo tanto, $T=215-141=74$. Final Answer: #### 74
Template 10 Example (English) Language: English Template: 10 Question: When Winnie watches her nephew, she gets out a variety of toys for him. The bag of building blocks has 72 blocks in it. The bin of stuffed animals has 47 stuffed animals inside. The tower of stacking rings has 29 multicolored rings on it. Winnie recently bought a tube of bouncy balls, bringing her total number of toys for her nephew up to 215. How many bouncy balls came in the tube? Answer: Let $T$ be the number of bouncy balls. Total number of toys before the balls: $72+47+29=148$. So $67148+T=215\Rightarrow T=215-148=$ 67. Final Answer: #### 67

### v2

@@ B.3 High to X Low-Resource Language Gaps
Aya-23-35B 51.21 57.4 54.12
Mixtral-8x7B-Instruct 43.66 56.8 57.11
Mistral-7B-Instruct 40.98 44.2 50.11
Aya-expanse-32B 56.87 71.6 60.45
Gemma-2-9b-it 52.29 64.4 52.44
Qwen3-8b 69.81 73.0 62.89
@@ Appendix D Full Functional Benchmark Results
AYA Exp-32B 74.93 55.52 49.05 46.90 49.86 18.06
AYA 23-35B 67.11 48.78 40.43 43.93 40.70 15.90
Mistral-7B-Instruct 59.57 41.78 40.97 25.33 21.83 18.59
Mixtral-8x7B-Instruct 56.06 36.39 33.70 31.53 24.65 12.40
Gemma-2-9B-it 75.20 53.36 50.40 47.16 48.24 22.91
Qwen3-8B 87.60 61.46 56.60 50.13 49.87 17.79
@@
English AYA 23-35B 67.11 73.75 70.89 77.41
Gemma2-9B-it 75.20 81.08 79.24 84.17
AYA Exp-32B 74.93 79.92 78.97 83.39
Mistral-7B 59.57 68.15 64.42 72.39
Mixtral-8x7B 56.06 64.48 61.19 69.50
Qwen3-8B 87.60 90.70 91.37 93.63
French AYA 23-35B 48.78 58.30 53.90 62.94
Gemma2-9B-it 53.36 61.38 56.06 63.89
AYA Exp-32B 55.52 63.12 58.22 65.83
Mistral-7B 41.78 51.35 46.90 56.18
Mixtral-8x7B 36.39 45.37 39.62 49.03
Qwen3-8B 61.46 69.11 63.07 70.46
Spanish AYA 23-35B 40.43 50.77 44.20 54.24
Gemma2-9B-it 50.40 59.07 51.75 61.00
AYA Exp-32B 49.05 58.49 51.21 60.62
Mistral-7B 40.97 50.96 43.66 53.86
Mixtral-8x7B 33.70 43.47 37.20 47.00
Qwen3-8B 56.60 65.06 57.41 65.83
Yoruba AYA 23-35B 15.90 22.77 16.17 23.74
Gemma2-9B-it 22.91 31.08 25.88 34.75
AYA Exp-32B 18.06 28.19 29.48 30.31
Mistral-7B 18.59 28.95 20.21 31.27
Mixtral-8x7B 12.40 20.27 15.00 23.36
Qwen3-8B 17.79 27.22 20.21 30.50
Arabic AYA 23-35B 43.93 53.47 46.90 56.37
Gemma2-9B-it 47.16 57.14 48.79 58.88
AYA Exp-32B 46.90 56.76 49.06 58.69
Mistral-7B 25.33 36.10 28.57 39.38
Mixtral-8x7B 31.53 40.73 35.85 45.95
Qwen3-8B 50.13 60.62 52.30 62.16
Hindi AYA 23-35B 40.70 50.19 44.47 54.44
Gemma2-9B-it 48.24 56.94 50.94 59.65
AYA Exp-32B 49.86 58.69 52.56 61.78
Mistral-7B 21.83 32.24 25.06 36.67
Mixtral-8x7B 24.65 33.24 32.83 41.31
Qwen3-8B 49.87 58.11 52.02 60.62
@@ F.2 CL-GSMSym Plots
Appendix G Translation Validation Rubrics
G.1 CL-IFEval Translation Validation (541 Prompts)
1. Objective The goal of this validation task is to ensure that every translated prompt accurately represents the intent, structure, and constraints of the original English version while remaining clear and natural in the target language. Annotators should: Confirm that the translation faithfully conveys the same meaning and instructions as the English source. Verify that every element in the instruction_id_list is valid and meaningful in the target language. Flag any instruction IDs that cannot be reasonably expressed or understood in that language. Correct minor translation errors or omissions directly in the prompt, but only when needed for consistency with the English version. Record every correction or flag for traceability.
2. Dataset Structure Each entry contains: key – unique identifier for the instance prompt – translated text (the only part that differs from English) instruction_id_list – the list of instructions applied to the English prompt kwargs – parameters or values associated with those instructions Example (simplified):
"key": 1001,
"prompt": "Translated prompt text...",
"instruction_id_list": ["punctuation:no_comma", "length_constraints:number_words"],
"kwargs": [{}, {"num_words": 300}]
3. Validation Criteria A. Translation Fidelity The translation must convey the same meaning, intent, and constraints as the English prompt. The tone and phrasing should match the formality and directive style of the English source. If the translation is missing, ambiguous, or distorts the meaning, revise the prompt to align with the English intent. When a fix is made, record it with a short note (e.g., “clarified no-comma rule,” “added minimum-count phrase”). Do not add new content or stylistic embellishments not present in the English prompt. B. Instruction–Prompt Alignment For every instruction in the instruction_id_list: * Realized – The translated prompt clearly expresses the instruction. * Not Applicable – The instruction refers to a feature or rule that cannot exist in the target language (e.g., case sensitivity in scripts that lack uppercase/lowercase). Keep the ID in the metadata, but flag it as “not applicable in this language.” Do not remove or modify IDs in the JSON. All flags are recorded separately. C. Linguistic and Logical Consistency The translation must be grammatically correct, culturally neutral, and logically coherent. Instructions like “Do not use commas” or “Write at least 300 words” must retain their restrictive force and avoid weaker formulations such as “Try not to use.” Quantitative and conditional phrases (e.g., “at least,” “exactly,” “no more than”) must match the English structure precisely. D. Formatting and Style Preserve examples, symbols, and URLs exactly as they appear. Maintain Markdown examples (*highlighted section*, headings, etc.) where relevant. Ensure punctuation, numbering, and quotation marks follow conventions appropriate for the target language. Do not introduce additional formatting not found in the original.
4. Correction and Flagging Process For each entry: 1. Compare the English prompt with the translated version. 2. Identify whether each instruction ID is not applicable. 3. If small translation inconsistencies exist, correct them directly in the prompt. 4. Record every modification or flag, even for minor fixes. 5. Do not introduce new instructions or stylistic changes that alter the intent. Permitted corrections include: Fixing missing or mistranslated phrases that change meaning. Clarifying numerical constraints or prohibitions (e.g., “Do not use commas”). Adjusting phrasing for grammatical correctness. Not permitted: Adding new examples, stylistic flourishes, or re-interpreting the prompt. Changing the number of constraints or rewriting the task.
5. Final Review and Consistency Check Before final submission: Ensure that all translations are accurate and natural in their respective languages. Verify that English-only or script-dependent instruction IDs (e.g., case-based rules) are consistently flagged as “Not Applicable.” Confirm that all modifications and flags have been logged. Export the validated, corrected dataset for each language along with its annotation log in the Log changes column.
G.2 CL-GSMSym Translations Validation Guidelines (100 Math Templates)
1. Objective The goal of this validation task is to ensure that every translated mathematical template accurately preserves the structure, logic, and functionality of the original English version. Annotators should verify that translations remain technically valid, mathematically consistent, and compatible with the generator framework. The task involves: Preserving all template markup ({…}, #init:, #conditions:, #answer:). Ensuring the math logic and units match the original. Correcting small formatting and spacing issues where necessary. Flagging every fix or inconsistency for audit tracking.
2. Dataset Components Each template file contains: question – the main text of the math problem. answer – the step-by-step explanation and final numeric answer. question_annotated – question with placeholders, variables, and meta blocks. answer_annotated – annotated solution including variables and computed forms. #init, #conditions, #answer – meta-sections used by the template generator. id_orig / id_shuffled – unique identifiers for the template.
3. Validation Criteria A. Template Markup Preservation Every variable token must remain identical to English: {var}, {var,label}, {x,10}, {k*y}, etc. The meta-sections (#init:, #conditions:, #answer:) must remain in English and unaltered. If a translation changed a placeholder (for example, {big_fish,shark} $\rightarrow$ {big_fish,requin}), restore the English version. Flag: “Restored placeholder token”. B. Math and Unit Consistency All numeric quantities, operations, and measurement units must match the logic in the English template. Example: if English uses “6 inches”, “10-foot shark”, “12 inches/foot”, the translation must also use imperial units, not metric. If the translation converted units (e.g. inches $\rightarrow$ cm, feet $\rightarrow$ meters), revert to the English units. Flag: “Unit drift fixed (cm $\rightarrow$ inches)”. C. Angle Brackets and Final Answer Maintain the <<...>> spans exactly as in the English version (these are computation traces). The final line of every answer must end with: #### <number> or #### {expression}. If extra spaces, broken symbols, or HTML entities appear, normalize and flag. Flag: “Computation span or final answer normalized”. D. Meta Sections (Technical Integrity) The #init:, #conditions:, and #answer: sections must remain in English and in the same order. If the translation system or annotator translated these lines, restore them verbatim from the original English. Flag: “Meta block reverted to EN”. Ensure indentation and syntax are unchanged — the generator relies on exact structure.
4. Correction and Flagging Process For each template: 1. Open both the English and translated versions. 2. Compare them line by line (especially question_annotated and answer_annotated). 3. Apply small corrections directly: Restore missing or broken placeholders. Fix spacing and punctuation. Revert units if changed. 4. Record every modification you make, even small ones. 5. If the template cannot be fixed to run logically, flag it for exclusion.
5. Final Review Checklist Before final submission, verify that: All placeholders {...} and computation spans <<...>> are intact. All templates end with the correct final answer format. Meta sections remain in English, properly indented. All numbers and units align with the original logic. Each modification has been recorded in your correction log.
6. Output The deliverables per language: A clean, generator-compatible set of 100 validated templates. A Column in the google sheet with every correction or flagged issue (for transparency and reproducibility).
Appendix H Failure Case Examples for CL-IFEval
Failure Example: length_constraints:number_words Model: Qwen3-8B Language: English Prompt: Create a riddle about the name Sheldon using only 10 words. Make sure to only use capital letters in your entire response. Model Response: SHELDON, A NAME WITH A FUNKY RING. Instructions: change_case:english_capital length_constraints:number_words Failure: The response contained only 8 words, violating the 10-word constraint.
Failure Example: change_case:capital_word_frequency Model: Qwen3-8B Language: French Prompt: Rédigez un CV pour Matthias Algiers. Utilisez des mots en majuscules pour mettre en avant les compétences clés, mais assurez-vous que les mots entièrement en majuscules apparaissent moins de 10 fois. Enveloppez toute la réponse entre guillemets. Model Response: "MATTHIAS ALGIERS RÉSUMÉ PROFESSIONNEL Professionnel dévoué et orienté résultats avec une solide expérience en GESTION DE PROJET, PLANIFICATION STRATÉGIQUE et LEADERSHIP D’ÉQUIPE…" Instructions: change_case:capital_word_frequency startend:quotation Failure: The response included more than 10 fully capitalized words, violating the capital word frequency constraint.
Appendix I Template Examples for CL-GSMSym
@@
Template 10 Example (Spanish) Language: Spanish Template: 10 Question: Cuando Emma observa a su primo, saca una variedad de juguetes para él. La bolsa de bloques de construcción tiene 74 bloques dentro. El contenedor de animales de peluche tiene 37 animales de peluche dentro. La torre de anillos apilables tiene 30 anillos multicolores. Emma compró recientemente un tubo de pelotas saltarinas, lo que eleva el número total de juguetes para su primo a 215. ¿Cuántas pelotas saltarinas vinieron en el tubo? Answer: Sea $T$ el número de pelotas que rebotan en el tubo. Después de comprar el tubo de pelotas, Emma tiene $74+37+30+T=141+T=215$ juguetes. Por lo tanto, T=215141=74T=215-141=74. Final Answer: #### 74
Template 10 Example (English) Language: English Template: 10 Question: When Winnie watches her nephew, she gets out a variety of toys for him. The bag of building blocks has 72 blocks in it. The bin of stuffed animals has 47 stuffed animals inside. The tower of stacking rings has 29 multicolored rings on it. Winnie recently bought a tube of bouncy balls, bringing her total number of toys for her nephew up to 215. How many bouncy balls came in the tube? Answer: Let $T$ be the number of bouncy balls. Total number of toys before the balls: $72+47+29=148$. So $148+T=215\Rightarrow T=215-148=67$. Final Answer: #### 67